A coffee seller has one grade of coffee selling for $4 per pound and another for $6.50 per pound. He mixes the coffee to sell at $4.75 per pound and mixes 150 pounds. How much of each type does he use?
How do you do this?

Answer by LilSkittleMd(119)

Let x=coffee-$4 per pound
Let y=coffee-$6.50 per pound
Set up the two equations
x+y=150
4x+6.50y=712.5
Use substitution method to solve.
x=150-y
4(150-y)+6.50y=712.5
600-4y+6.5y=712.5
Combine like terms
600+2.5y=712.5
2.5y=112.5
y=45
Now substitute y into one of the original equations
x+45=150
x=105
105 pounds of the $4 per pound coffee was used
45 pounds of the $6.50 per pound coffee was used